._____._________._____.
A B C D
Statements Reasons
AC=BD GIVEN
AC=AB+BC Def. of Betweenness
BD= BC+CD Def. of Betweenness
AB+BC=BC+CD SUBSTITUTION PROPERTY
BC=BC REFLEXIVE PROPERTY
AB=CD TRANSITIVE PROPERTY

Found 3 solutions by midwood_trail, deanice_01@yahoo.com, smartguy101:
Answer by midwood_trail(310)

The way you typed this question makes no sense. This is why no one has been able to help you with this proof.
What exactly are you trying to prove here?
Can you retype it?

Answer by deanice_01@yahoo.com(3)

._____._________._____.
A B C D
Statements Reasons
AC=BD GIVEN
AC=AB+BC Def. of Betweenness
BD= BC+CD Def. of Betweenness
AB+BC=BC+CD SUBSTITUTION PROPERTY
BC=BC REFLEXIVE PROPERTY
AB=CD TRANSITIVE PROPERTY


Answer by smartguy101(5)

deanice, u r really really really really really really really really really really really really really really really really HOT!!!!!!!!!!!!!!!!!!!!!!!!!!!1